OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of square root of the quantity 16 r to the ninth power. The total number of tiles used can be represented by the expression below. one hundred twenty r to the third power, all over the square root of the quantity sixteen r to the ninth power

OpenStudy (anonymous):

@ankit042 i need more help pleasee

OpenStudy (ankit042):

try solving the denominator first

OpenStudy (anonymous):

sqrt(8r^3)?

OpenStudy (ankit042):

NO it will be 4r^3*sqrt(r^3)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

is that a 9/3?

OpenStudy (anonymous):

@ybarrap why did you delete that :( i wanted you to explain it to me

OpenStudy (ankit042):

I think it will be 4r^3*sqrt(r^3) like the last time we did 16r^9 can be written as 4r^3*4r^3*r^3 after putting sqrt we will get 4r^3*sqrt(r^3)

OpenStudy (ybarrap):

$$ \tt \large\frac{120r^3}{\sqrt{16r^9}} \large=\frac{120r^3}{\sqrt{16}r^{9/2}} \large=\frac{120r^{3-9/2}}{4}\\ \large=\frac{120r^{-3/2}}{4} \large=\frac{120}{4r^{3/2}}\\ \large=\frac{30}{r^{3/2}}\\ $$

OpenStudy (ybarrap):

let me know if you have any questions

OpenStudy (anonymous):

im just looking over it and seeing what you did step by step (: